# Discrete Chi-Square Method can model and forecast complex time series

### v5

@@ Abstract
We show how intensive, large and accurate time series can allow us to see through time. Many phenomena have aperiodic and periodic components. An ideal time series analysis method would detect such trend and signal(-s) combinations. The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations constraining the trend and signal(-s) detection. We show that none of those limitations constrains our Discrete Chi-square Method (DCM) which can detect signal(-s) superimposed on an unknown trend. Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency. The DCM’s backbone is the Gauß-Markov theorem that the Least Squares (LS) is the best unbiased estimator for linear regression models. DCM can not fail because this simple method is based on the computation of a massive number of linear model LS fits. The Fisher-test gives the signal significance estimates and identifies the best DCM model from all alternative tested DCM models. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational well-posed solution. The DCM can forecast complex time series. The best DCM model must be correct if it passes our Forecast-test. Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series. In our appendix, we show that DCM can model and forecast El Niño.
@@
Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform, El Niño
@@ 1 Introduction
This study proceeds through following stages. We present the DCM formulation (Section 2.1). The WD-effect is discussed (Section 2.2). Complex time series are simulated using the DCM model (Sect. 2.3). The DFT and its ALs are presented (Section 2.4). We compare our DCM to the renowned DFT. The time series analysis of simulated data sets shows that the ALs of DFT do not constrain DCM (Sections 3.1-3.8). Then, we present the identification of the best DCM model for the data (Section 3.9), the DCM significance estimates (Section 3.10), the well-posed computational DCM model solution (Section 3.11) and the DCM forecasts (Section 3.12). Finally, we discuss our results (Section 4).
@@ 2.1 Discrete Chi-square method (DCM)
There are, of course, totally wrong DCM models for the data. For example, DCM can be forced search for too few or too many $K_{1}$ signals, or the selected $p(t)$ trend order $K_{3}$ can be wrong, as shown in Figures 5-10 by [27]. Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in Jetsu [^1] Jetsu L. “Do planets cause the sunspot cycle?”, submitted to Scientific Reports. August 13th, 2025. (2025). These unstable models have three signatures
@@
The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [29, TSPA, Equation 1] and the Continuous Period Search [36, CPS, Equation 3]. TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$. TSPA can detect pure sine and double wave signals ($K_{2}=1$ or 2). The extension of TSPA, the CPS method, tests three alternatives: $K_{2}=1$ or 2, or no signal at all. The DCM model is more sophisticated because it can have any arbitrary $K_{1}$, $K_{2}$ and $K_{3}$ combination. The DCM sum $g(t)$ =h(t)+p(t)=h(t)+p(t) of arbitrary periodic and aperiodic functions represents a universal model because the innumerable $K_{1}$, $K_{2}$ and $K_{3}$ combinations allow unlimited complexity. Even more complex signal shape $(K_{2})$ combinations could be used. For example, these shapes could be $K_{2,i=1}=3$ (3rd harmonic), $K_{2,i=2}=1$ (pure sine) and $K_{2,i=3}=2$ (double wave), where if $K_{2,i}$ denotes the shape of $i$:th signal.
@@ 2.4 Discrete Fourier Transform (DFT)
The DFT is one of the most prevalent frequency-domain parametric time series analysis methods for unevenly spaced time series. It searches for the best pure sine model for the data. The equivalent DCM model has the orders $K_{1}=1,K_{2}=1$ and $K_{3}=0$. Any time series analysis method must be remarkable if it performs better than the distinguished DFT. Therefore, we will compare how DCM and DFT perform in the analyses of simulated time series (Sections 3.1-3.7). We search for signals in these simulated time series by using the frequently applied DFT version formulated by [23] [^2] This paper had over 2600 citations in December 2025. Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [23, Equation 1]. The notations for the DFT model are
@@
The time series is “too short” [30, AL6]
@@
The time series is non-stationary because it contains “a trend” [42, AL7]
@@
Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [30, AL9].
@@ 3.1 Model 1
Model 1 is a DCM model, where $K_{1}=1$, $K_{2}=1$, $K_{3}=0$, $B_{1,1}=$ $(A_{1}/2)\cos{(2\pi f_{1}t_{\mathrm{1,max,1}})}$ and $C_{1,1}=$ $(A_{1}/2)$ $\sin{(2\pi f_{1}t_{\mathrm{1,max,1}})}.$ We give the DCM analysis results for three samples having different $n$ and $\mathrm{SN}$} combinations (Table 1: Columns 2-4). For each sample, this table specifies the electronic data file and the electronic DCM control file. [^3] We publish all our data files and all DCM control files in https://zenodo.org/uploads/17018676 The detected $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values are correct and accurate even for the combination $n=50$ and ${\mathrm{SN}}=50$. Regardless of $\Delta T<P_{1}$, these model parameter values become more accurate and converge to the correct simulated values when $n$ and $\mathrm{SN}$} increase. This confirms the WD-effect.
@@ 3.3 Model 3
The DFT detects the wrong period $P_{1}=1.163$ for the original data (Figure 3e: diamond). This is an expected result because the detected period should be close to $(P_{1}+P_{2})/2$ when the peak to peak amplitudes of the simulated data, $A_{1}=A_{2}$, are equal (Jetsu 2025). The two DFT periodogram $z_{\mathrm{DFT}}(f)$ peaks at frequencies $1/P_{1}$ and $1/P_{2}$ overlap and merge into one peak. The period $P_{2}=0.182$ detected for the residuals is also wrong (Figure 3f: diamond). The DFT trend $p_{\mathrm{PDF}}(t)$ estimate $M_{0}=0.880$ fails (Figure 3f: dashed black line). The black dots $y_{i}$ show minor deviations from the continuous black DFT model $g_{\mathrm{DFT}}(t)$ line (Figure 3g). The detrended model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (continuous black line), the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots), the pure sine signal $s_{\mathrm{y,DFT}}(t)$ for the data (continuous thick red line) and the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ for the residuals (continuous blue thin line) are shown in Figure 3h. Note that the $s_{\mathrm{y,DFT}}(t)$ and $s_{\mathrm{\epsilon,DFT}}(t)$ signal amplitudes are far from equal. The DFT residuals (blue dots) offset to the level of -3.0 (blue dotted line) are not stable.
@@ 3.4 Model 4
Since the peak to peak amplitudes of the simulated data signals are equal, $A_{1}=A_{2}$, the expected result for the DFT analysis of original data is $(P_{1}+P_{2})/2$, where $P_{1}$ and $P_{2}$ are the simulated signal periods (Jetsu 2025). The DFT detects this expected wrong period $P_{1}=1.165$ for the original data (Figure 4e: diamond). A wrong period $P_{2}=0.147$ is also detected for the residuals (Figure 4f: diamond). The DFT estimate $M_{0}=0.962$ for the $p(t)$ trend is close to the correct value $M_{0}=1$, but the $M_{1}=-0.017$ and $M_{2}=0.903$ estimates are wrong (Figure 4g: dashed black line). The continuous black DFT model $g_{\mathrm{DFT}}(t)$ line deviates from the black dots of data $y_{i}$, especially in the beginning and end of the sample. In our Figure 4h, the black dots are the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ and the continuous black line is the detrended DFT model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$. The continuous thick red line is the pure sine signal $s_{\mathrm{y,DFT}}(t)$ for the original data and the continuous thinner blue line is the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ for the residuals. The $s_{\mathrm{\epsilon,DFT}}(t)$ signal amplitude is far below the simulated correct value $A_{2}=2.0$. The blue dots representing the DFT model residuals are offset to the level of -3.0 and show clear trends.
@@ 3.8 Summary of simulations
Our DCM outperforms the esteemed DTF. The DCM detects the correct signal(s) and trend, but DFT does not. For the sunspot data, the DCM detects signals superimposed on a constant trend (Jetsu 2025). The simulations presented here support the conclusion that those solar signals and the trend are correct. This would explain why the DCM is the first time series analysis method that detects Jupiter’s exact $11.^{\mathrm{y}}86$ period in the sunspot record. The one-dimensional time series analysis methods, like DFF, have failed to find exactly correct periods.
Table 8: Fisher tests. DCM analysis of Model 1 simulated time series ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100). (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name. (2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.33footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
@@ 3.9 Best model
In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$. We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large. One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu (2025, Table S7).
@@ 3.10 Significance estimates
Jetsu [27] or Jetsu (2025) gave no signal significance estimate for the first detected period. Here, we use the Fisher-test for this purpose. The one signal model is the complex model. The logical simple model alternative is $g(t)=m\equiv$ white noise having standard deviation $s$. However, white noise is not the only possible alternative simple model. The other possible nested no signal DCM polynomial models are
@@
After the detection of the first strongest signal, the Fisher-test critical level $Q_{\mathrm{F}}$ values increase for the next detected weaker signals. In other words, the signal significances decrease. Typical examples can be found in Jetsu 2025 (Tables S5-S16). No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected. For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
For this Model 1 time series, there is no need to discuss the DFT significance estimates [23, their Equation 22] because this method fails to detect the correct period.
@@
Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated time series (electronic data file Model1n100SN100). (1) Polynomial model (Equation 49). (2) DCM model $g_{1,1,0}$. Note that $\mathcal{M}$=1-3  polynomials represent simple models, and $\mathcal{M}$=5-8  polynomials represent complex models. Otherwise as in Table 8. 44footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.

### v6

@@ Abstract
Forecasting El Niño is one of the greatest challenges of science. We show how intensive, large and accurate time series allow us to see through time. Our Discrete Chi-square Method (DCM) can detect arbitrary trend and signal(-s) combinations. It can forecast complex time series. The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations. None of those limitations constrains the DCM. Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency. The DCM’s backbone is the Gauß-Markov theorem that the Least Squares (LS) is the best unbiased estimator for linear regression models. DCM can not fail because this simple method is based on the computation of a massive number of linear model LS fits. The Fisher-test gives the signal significance estimates and identifies the best DCM model from all alternative tested DCM models. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational well-posed solution. The best DCM model must be correct if it passes our Forecast-test. Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series. In our appendix, we show that the DCM can model and forecast El Nino data between 1870 and 2024. An immediate, independent and objective validity check of our analysis may save some money.
@@
El Niño, Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform
@@ 1 Introduction
This study proceeds through following stages. We present the DCM formulation (Section 2.1). The WD-effect is discussed (Section 2.2). Complex time series are simulated using the DCM model (Sect. 2.3). The DFT and its ALs are presented (Section 2.4). We compare our DCM to the renowned DFT. The time series analysis of simulated data sets shows that the ALs of DFT do not constrain DCM (Sections 3.1-3.8). Then, we present the identification of the best DCM model for the data (Section 3.9), the DCM significance estimates (Section 3.10), the well-posed computational DCM model solution (Section 3.11) and the DCM forecasts (Section 3.12). Finally, we discuss our results (Section 4). In our Appendix, we show that the DCM can model and forecast El Niño (Section A1).
@@ 2.1 Discrete Chi-square method (DCM)
There are, of course, totally wrong DCM models for the data. For example, DCM can be forced search for too few or too many $K_{1}$ signals, or the selected $p(t)$ trend order $K_{3}$ can be wrong, as shown in Figures 5-10 by [29]. Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in [30]. These unstable models have three signatures
@@
The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [31, TSPA, Equation 1] and the Continuous Period Search [39, CPS, Equation 3]. TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$. TSPA can detect pure sine and double wave signals ($K_{2}=1$ or 2). The extension of TSPA, the CPS method, tests three alternatives: $K_{2}=1$ or 2, or no signal at all. The DCM model is more sophisticated because it can have any arbitrary $K_{1}$, $K_{2}$ and $K_{3}$ combination. The DCM sum $g(t)$ =h(t)+p(t)=h(t)+p(t) of arbitrary periodic and aperiodic functions represents a universal model because the innumerable $K_{1}$, $K_{2}$ and $K_{3}$ combinations allow unlimited complexity. Even more complex signal shape $(K_{2})$ combinations could be used. For example, these shapes could be $K_{2,i=1}=3$ (3rd harmonic), $K_{2,i=2}=1$ (pure sine) and $K_{2,i=3}=2$ (double wave), where if $K_{2,i}$ denotes the shape of $i$:th signal.
@@ 2.4 Discrete Fourier Transform (DFT)
The DFT is one of the most prevalent frequency-domain parametric time series analysis methods for unevenly spaced time series. It searches for the best pure sine model for the data. The equivalent DCM model has the orders $K_{1}=1,K_{2}=1$ and $K_{3}=0$. Any time series analysis method must be remarkable if it performs better than the distinguished DFT. Therefore, we will compare how DCM and DFT perform in the analyses of simulated time series (Sections 3.1-3.7). We search for signals in these simulated time series by using the frequently applied DFT version formulated by [24] [^1] This paper had over 2600 citations in December 2025. Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [24, Equation 1]. The notations for the DFT model are
@@
The time series is “too short” [32, AL6]
@@
The time series is non-stationary because it contains “a trend” [45, AL7]
@@
Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [32, AL9].
@@ 3.1 Model 1
Model 1 is a DCM model, where $K_{1}=1$, $K_{2}=1$, $K_{3}=0$, $B_{1,1}=$ $(A_{1}/2)\cos{(2\pi f_{1}t_{\mathrm{1,max,1}})}$ and $C_{1,1}=$ $(A_{1}/2)$ $\sin{(2\pi f_{1}t_{\mathrm{1,max,1}})}.$ We give the DCM analysis results for three samples having different $n$ and $\mathrm{SN}$} combinations (Table 1: Columns 2-4). For each sample, this table specifies the electronic data file and the electronic DCM control file. [^2] We publish all our data files and all DCM control files in
https://zenodo.org/records/18726181/
The detected $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values are correct and accurate even for the combination $n=50$ and ${\mathrm{SN}}=50$. Regardless of $\Delta T<P_{1}$, these model parameter values become more accurate and converge to the correct simulated values when $n$ and $\mathrm{SN}$} increase. This confirms the WD-effect.
@@ 3.3 Model 3
The DFT detects the wrong period $P_{1}=1.163$ for the original data (Figure 3e: diamond). This is an expected result because the detected period should be close to $(P_{1}+P_{2})/2$ when the peak to peak amplitudes of the simulated data, $A_{1}=A_{2}$, are equal [30]. The two DFT periodogram $z_{\mathrm{DFT}}(f)$ peaks at frequencies $1/P_{1}$ and $1/P_{2}$ overlap and merge into one peak. The period $P_{2}=0.182$ detected for the residuals is also wrong (Figure 3f: diamond). The DFT trend $p_{\mathrm{PDF}}(t)$ estimate $M_{0}=0.880$ fails (Figure 3f: dashed black line). The black dots $y_{i}$ show minor deviations from the continuous black DFT model $g_{\mathrm{DFT}}(t)$ line (Figure 3g). The detrended model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (continuous black line), the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots), the pure sine signal $s_{\mathrm{y,DFT}}(t)$ for the data (continuous thick red line) and the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ for the residuals (continuous blue thin line) are shown in Figure 3h. Note that the $s_{\mathrm{y,DFT}}(t)$ and $s_{\mathrm{\epsilon,DFT}}(t)$ signal amplitudes are far from equal. The DFT residuals (blue dots) offset to the level of -3.0 (blue dotted line) are not stable.
@@ 3.4 Model 4
Since the peak to peak amplitudes of the simulated data signals are equal, $A_{1}=A_{2}$, the expected result for the DFT analysis of original data is $(P_{1}+P_{2})/2$, where $P_{1}$ and $P_{2}$ are the simulated signal periods [30]. The DFT detects this expected wrong period $P_{1}=1.165$ for the original data (Figure 4e: diamond). A wrong period $P_{2}=0.147$ is also detected for the residuals (Figure 4f: diamond). The DFT estimate $M_{0}=0.962$ for the $p(t)$ trend is close to the correct value $M_{0}=1$, but the $M_{1}=-0.017$ and $M_{2}=0.903$ estimates are wrong (Figure 4g: dashed black line). The continuous black DFT model $g_{\mathrm{DFT}}(t)$ line deviates from the black dots of data $y_{i}$, especially in the beginning and end of the sample. In our Figure 4h, the black dots are the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ and the continuous black line is the detrended DFT model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$. The continuous thick red line is the pure sine signal $s_{\mathrm{y,DFT}}(t)$ for the original data and the continuous thinner blue line is the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ for the residuals. The $s_{\mathrm{\epsilon,DFT}}(t)$ signal amplitude is far below the simulated correct value $A_{2}=2.0$. The blue dots representing the DFT model residuals are offset to the level of -3.0 and show clear trends.
@@ 3.8 Summary of simulations
Our DCM outperforms the esteemed DTF. The DCM detects the correct signal(s) and trend, but DFT does not. For the sunspot data, the DCM detects signals superimposed on a constant trend [30]. The simulations presented here support the conclusion that those solar signals and the trend are correct. This would explain why the DCM is the first time series analysis method that detects Jupiter’s exact $11.^{\mathrm{y}}86$ period in the sunspot record. The one-dimensional time series analysis methods, like DFF, have failed to find exactly correct periods.
Table 8: Fisher tests. DCM analysis of Model 1 simulated time series ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100). (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name. (2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
@@ 3.9 Best model
In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$. We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large. One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu [30, Table S7].
@@ 3.10 Significance estimates
Jetsu [29] or Jetsu [30] gave no signal significance estimate for the first detected period. Here, we use the Fisher-test for this purpose. The one signal model is the complex model. The logical simple model alternative is $g(t)=m\equiv$ white noise having standard deviation $s$. However, white noise is not the only possible alternative simple model. The other possible nested no signal DCM polynomial models are
@@
After the detection of the first strongest signal, the Fisher-test critical level $Q_{\mathrm{F}}$ values increase for the next detected weaker signals. In other words, the signal significances decrease. Typical examples can be found in Jetsu [30, Tables S5-S1]. No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected. For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
For this Model 1 time series, there is no need to discuss the DFT significance estimates [24, their Equation 22] because this method fails to detect the correct period.
@@
Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated time series (electronic data file Model1n100SN100). (1) Polynomial model (Equation 49). (2) DCM model $g_{1,1,0}$. Note that $\mathcal{M}$=1-3  polynomials represent simple models, and $\mathcal{M}$=5-8  polynomials represent complex models. Otherwise as in Table 8. 33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.